### v5

@@ I Introductory Remarks
While 3GPPprovides foundational support for RAN slicing, the O-RAN Alliance plays a complementary and critical role in enabling intelligent slicing through open interfaces and components [43, 23]. To unlock the full potential of slicing within O-RANframework, a strong unification between O-RAN and 3GPP standards is essential. It helps to address the complex multi-domain orchestration challenges associated with RAN slicing [44]. Beyond traditional RAN slicing, O-RAN introduces several key enhancements that improve customization, interoperability and automation: (a) Through its disaggregated architecture, O-RAN enables fine-grained control over slice specific resources. Operators can select specialized hardware and software from different vendors to optimize the performance per slice. (b) O-RAN’s open interfaces eliminates vendor lock-in, making it possible to mix and match interoperable components. This allows operators to choose the best-in-breed solutions for each slice, leading to cost savings and innovation. (c) The integration of non-real-time RIC (Non-RT RIC) and near-real-time RIC (Near-RT RIC) introduces intelligent control via Non-RT RIC applications (rApps) and Near-RT RIC applications (xApps), which enable close-loop automation and optimization of slice-specific resources management [45, 46, 47]. (d) O-RANfurther supports automation of slice lifecycle management–encompassing slice creation, configuration, and orchestration–through its interfaces and protocols, significantly reducing operational complexity and improving reliability.
Note that, while both O-RAN and 3GPPaddresses key aspects of RAN slicing, neither framework explicitly focuses on the virtualization aspects. Thus, a broader harmonization—encompassing O-RAN, 3GPP, and European Telecommunications Standards Institute (ETSI)—is required to support both the physical and virtual infrastructure. Such alignment is anticipated to deliver unified and scalable solutions for the realization of flexible, service-specific O-RAN slices [44].
@@ II-C O-RAN Standards and Activities
The introduction of functional split in 3GPP represents a critical step towards disaggregating the standard protocol stack. This involves separating the processing of a specific layer within the protocol stack from the computing entity, thereby promoting architectural openness, intelligent cellular interface, and the feasibility of network slicing [52]. This functional split has served as a pivotal catalyst for the development of subsequent O-RAN-related specifications, which will be elaborated upon later in this section.
@@ II-C3 Telecom Infra Project
They concentrate on developing integrated RAN solutions tailored to specific deployment scenarios across varied environments, including both indoor and outdoor use cases.
@@ II-C4 Small Cell Forum
The small cell forum (SCF) is a global organization focused on developing technical specifications and commercial tools to accelerate the adoption of flexible, cost-effective, and scalable cellular network infrastructure. Throughout its existence, SCF has played a key role in standardizing essential elements of mobile network technology, including the functional API (FAPI), network FAPI (nFAPI), and enhancement to the X2 interface. These specifications enable an open, multi-vendor platform, thereby reducing the barriers to the densification of stakeholders in the wireless communications industry [98].
@@ II-D1 ONAP
ONAPcollaborates closely with the OSC, particularly on deploying the SMO and integrating the Non-RT RIC functionalities [72]. This partnership enhances coordination, minimizes duplication of efforts, and streamlines development. Shared priorities are outlined in [100], including a joint study on the “ONAP/3GPP & O-RAN Alignment-Standards Defined Notifications over VES” use case, which seeks to align ONAP with O-RAN Alliance and 3GPP, fostering interoperability and broader adoption [101]. The Kohn release of ONAPfurther advances this integration by enhancing cloud-native NF orchestration, supporting intent-driven closed-loop automation, and E2E network slicing across 5G RAN, CN, and TN domains [102, 103].
@@ II-D2 OpenAirInterface
In the context of 5G, the OAI community and software assets have been expanding rapidly. Current active projects include: OAI 5G RAN, OAI 5G CORE, MOSAIC 5G (M5G), and continuous integration (CI)/continuous deployment (CD). The newly created M5G PG aims to transform both the RAN and CN into agile and open platform for network service delivery. The M5G PGfocuses on developing software implementations of the O-RAN E2 protocol, as well as FlexRIC (a flexible RIC), FlexCN (a flexible core controller), and intelligent orchestration tools for RAN and CN domains [104]. Additionally, researchers at Northeastern University have successfully integrated OAI with the OSC RIC, enabling interoperability between open source development and standardized O-RAN implementations. Further details are provided in [105].
@@ III-B9 NG Interface
The NG interface, adopted by O-RAN from 3GPP, connects the O-CU to the 5G core (5GC). It comprises two components: NG-C for the control plane and NG-U for the user planes. NG-C connects the O-CU-CP with the access and mobility management function (AMF), while NG-U connects the O-CU-UP to the user plane function (UPF) in 5GC [20].
@@ III-C RAN Intelligent Controller
Moreover, the RIC is responsible for configuring network slices, orchestrating NFs, monitoring network performance, and conducting real-time optimizations of RAN resources through standardized open interfaces [147, 149]. As illustrated in Figure 2, the RICis implemented in two distinct forms—Non-RT RIC and Near-RT RIC—each designed to address specific control loop dynamics and latency requirements [138]. The following sections provide a detailed examination of both variants, elaborating on their functionalities, applications, and strategic importance in the broader context of RAN optimization and automation.
@@ III-D The SMO Framework and its Corresponding Interfaces
The architecture of SMO—particularly the Non-RT RIC—offers flexible implementation options, enabling the operators to selectively adopt the desired features. The SMO can also integrate with an E2E multi-domain service orchestrator, coordinating domain-specific modules across the RAN, TN, and CN. This E2Eorchestration facilitates on-demand creation and E2E management of network slice instances over distributed 5G infrastructure [166].
@@ III-D2 O2 Interface
In addition to lifecycle management of O-Cloud infrastructure, the O2 interface facilitates the orchestration of resource management, such as inventory, monitoring, provisioning, software management. It provide logical services that governs the deployment of O-RAN NFs using cloud resources. Although the Non-RT RIC may be deployed alongside other SMO components, the lifecycle is not expected to be managed through the O2 interface.
The O2 interface is inherently extensible, allowing new data types or features to be incorporated without altering the underlying protocol or management processes. It is also vendor-neutral and is unaffected by particular SMO framework and O-Cloud implementations. Through the O1 and/or O2 interfaces, operators can dynamically manage, reconfigure, and upgrade network components hosted within the O-Cloud environment [169].
@@
The 3GPP-defined MFs for the M&O of network slicing include the communication service management function (CSMF), network slice management function (NSMF), network slice subnet management function (NSSMF), and network function management function (NFMF) [38]. These MFs can be employed within the SMO framework, in accordance with the requirements outlined by the 3GPP specifications. In addition, the O-RAN Alliance has defined several functional and non-functional requirements for network slicing architecture as detailed in [45]. The requirements specified by both the 3GPP and the O-RAN Alliance are essential for the successful implementation of network slicing within the O-RAN architecture and for ensuring the effective operation of the 3GPP MFs. The provision of management service (MnS) for a mobile network, which includes network slicing, can be achieved through a set of functional blocks, as illustrated in Figure 3.
@@
In the following, we provide a detailed explanation of the previously mentioned 3GPP MFs and their role in the realization of the M&O of O-RAN network slice instances.
@@ III-E3 O-RAN Cloud Platform
The O-RAN cloud platform (O-Cloud)refers to a cloud computing environment composed of physical infrastructure compatible with O-RAN. It supports the deployment of critical O-RAN components such as the SMO, Near-RT RIC, and E2 nodes, along with the associated software and required M&O services. Each O-Cloud is composed of a group of central processing units (CPUs), random access memory (RAM), storage, network interface cards (NICs), basic input and output system (BIOS), baseboard management controllers (BMCs), and hardware accelerators. These elements work together to handle computationally intensive tasks across the platform [20].
@@ IV-B3 Service Layer
It is important to recognize that while the requirements for PNFs, cloudified network services, and O-Cloud platforms may vary depending on the situation. However, the logical network function requirements always remain the same [45, 168]. For Example, in the scenario illustrated, a single O-CU-CP instance manages the control plane operations for both network slices, whereas each slice is assigned a dedicated O-CU-UP instance. If the UE is connected to both slices, only one RRC connection is established, handling handover and cell assignments through the shared O-CU-CP. However, each service associated with a different network slice instances (NSI) can benefits personalized QoS management and independent flow control through an individual SDAP/PDCP stacks within their respective O-CU-UP [45].
@@ IV-C SMO Framework Deployment Options
The EXT-API offers northbound interoperability for the ONAP platform, serving as an access point for third-party frameworks. Upon receiving a service request, EXT-API produces a Service Order ID and transmits it in response, which can be used to track and monitor the status of the service order. Subsequently, the EXT-API activates the service orchestrator (SO) API to initiate the actual service creation process. This action represents progress in establishing uniform external interfaces for automated network slice orchestration [103, 188].
@@ V-C Xhaul Transport Network
V-C1 Xhaul Transport Network Architecture
@@
Underlay networks form the physical infrastructure of a TN. They comprises Ethernet switches, routers, dense WDM (DWDM) equipments, and the fiber optic cabling that interconnects these components into a network topology. To support an Xhaul TN environment, the packet switched network must handle both L2 and L3 services. Currently, L2 underlay networks predominantly rely on Ethernet, often utilizing virtual LANs (VLANs) for segmentation. Within O-RAN, following two prevalent packet switched underlay technologies are discussed in [192].
@@
V-C2 Xhaul Transport Network Slicing
The packet switched TNs provide a robust framework to support the network slicing. The Xhaul TN infrastructure is designed to support the diverse transport demands of various interfaces. They require tailored solutions for control, management, and user plane interfaces. This includes 5G TN segments, where FH, MH, and BH interfaces require tailored resource allocation [77, 201]. Each interface has specific latency, bandwidth, and traffic demands, which slicing helps to address by enabling precise resource customization [202].
@@ VI-B2 Slow Loop RAN Slice SLA optimization
In the second option the SMO framework updates the slice configuration of the Near-RT RIC or the RAN nodes based on instructions from the AI/ML model or control app. After the update, two outcomes are possible: either the Near-RT RIC and the RAN nodes implement the updated slice configuration, or the Near-RT RIC receives the updated A1 policies, take control of the RAN nodes and provide feedback to the Non-RT RIC via the A1 interface.
@@ VII-A Lessons Learned Related to the Architecture of O-RAN
We discussed on several occassions that the O-RAN architecture is composed of key components and interfaces that support its open, intelligent, and modular design. We gained several valuable insights from our study on these components and interfaces. The key takeaways are summarized below.
Fully Disaggregated Cloud-based RAN: The adoption of O-RAN marks a pivotal shift towards fully disaggregated RAN architectures, where key O-gNB functions are modularized into distinct components such as the O-CU, O-DU, and O-RU. The key contribution of O-RAN Alliance is the definition of O-FH interface that splits the DU into O-DU and O-RU. We learned that such a disaggregation can play a significant role in the isolation of O-RAN slices. To further enhance isolation, resource efficientcy, and cloud-native solutions, we beleive that it would of interested of network operators and vendors to further split the O-CU and O-DU into micro VNFs, such as virtual network function components (VNFCs) defined by the ETSI ISG NFV.
@@ IX Concluding Remarks and Future Outlook
In conclusion, the exploration of O-RAN illuminates its transformative potential within the context of cellular communications systems. As the wireless industry evolves to meet the demands of 5G, 6G, and beyond cellular systems, O-RAN emerges as a promising paradigm shift, offering flexibility, interoperability, and cost-efficiency in telecommunications networks deployment and management. Through our comprehensive analysis in this survey and tutorial paper, it becomes evident that O-RAN’s disaggregated approach to wireless network elements, enabled by open interfaces, automation, intelligence, and software-defined networking principles, fosters innovation and competition among vendors while reducing vendor lock-in. This approach not only spurs the development of diverse and specialized network functions but also empowers operators to tailor their networks to specific use cases and environments with greater agility and granularity through the deployment of network slicing at both network and management domains. In essence, while the journey towards realizing the full potential of network slicing in O-RAN may be fraught with several challenges, the destination promises a network architecture that is more open, agile, and responsive to the evolving needs of wireless communication in the next decade.

### v6

@@ I Introductory Remarks
While 3GPPprovides foundational support for RAN slicing, the O-RAN Alliance plays a complementary and critical role in enabling intelligent slicing through open interfaces and components [43, 23]. To unlock the full potential of slicing within O-RANframework, a strong unification between O-RAN and 3GPP standards is essential. Beyond traditional RAN slicing, O-RAN introduces several key enhancements that improve customization, interoperability and automation: (a) Through its disaggregated architecture, O-RAN enables fine-grained control over slice specific resources. Operators can select specialized hardware and software from different vendors to optimize the performance per slice. (b) O-RAN’s open interfaces eliminates vendor lock-in, making it possible to mix and match interoperable components. This allows operators to choose the best-in-breed solutions for each slice, leading to cost savings and innovation. (c) The integration of non-real-time RIC (Non-RT RIC) and near-real-time RIC (Near-RT RIC) introduces intelligent control via Non-RT RIC applications (rApps) and Near-RT RIC applications (xApps), which enable close-loop automation and optimization of slice-specific resource management [44, 45, 46]. (d) O-RAN further supports automation of slice lifecycle management–encompassing slice creation, configuration, and orchestration–through its interfaces and protocols, significantly reducing operational complexity and improving reliability.
Note that, while both O-RAN and 3GPP address key aspects of RAN slicing, neither framework explicitly focuses on the virtualization aspects. Thus, a broader harmonization—encompassing O-RAN, 3GPP, and European Telecommunications Standards Institute (ETSI)—is required to support both the physical and virtual infrastructure. Such alignment is anticipated to deliver unified and scalable solutions for the realization of flexible, service-specific O-RAN slices [47].
@@ II-C O-RAN Standards and Activities
The introduction of functional split in 3GPP represents a critical step towards disaggregating the standard protocol stack. This involves separating the processing of a specific layer within the protocol stack from the computing entity, thereby promoting architectural openness, intelligent cellular interface, and the feasibility of network slicing [52]. This functional split has served as a pivotal catalyst for the development of subsequent O-RAN-related specifications.
@@ II-C3 Telecom Infra Project
They concentrate on developing integrated RAN solutions tailored to diverse deployment scenarios across indoor and outdoor use cases.
@@ II-C4 Small Cell Forum
The small cell forum (SCF) is a global organization focused on developing technical specifications and tools to accelerate the adoption of flexible, cost-effective, and scalable cellular network infrastructure. The SCF has played a key role in standardizing essential elements of network technology, including the functional API (FAPI), network FAPI (nFAPI), and enhancement to the X2 interface. These specifications enable an open, multi-vendor platform, thereby reducing the barriers to the densification of stakeholders in the wireless communications industry [98].
@@ II-D1 ONAP
ONAPcollaborates with the OSC, particularly on deploying the SMO and integrating the Non-RT RIC functionalities [72]. This partnership enhances coordination, minimizes duplication of efforts, and streamlines development. Shared priorities are outlined in [100], including a joint study on the “ONAP/3GPP & O-RAN Alignment-Standards Defined Notifications over VES” use case, which seeks to align ONAP with O-RAN Alliance and 3GPP, fostering interoperability and broader adoption [101]. The Kohn release of ONAP further advances this integration by enhancing cloud-native NF orchestration, supporting intent-driven closed-loop automation, and E2E network slicing across 5G RAN, CN, and TN domains [102, 103].
@@ II-D2 OpenAirInterface
In the context of 5G, the OAI community and software assets have been expanding rapidly. Current active projects include: OAI 5G RAN, OAI 5G CORE, MOSAIC 5G (M5G), and continuous integration (CI)/continuous deployment (CD). The newly created M5G PG aims to transform both the RAN and CN into agile and open platform for network service delivery. The M5G PGfocuses on developing software implementations of the O-RAN E2 protocol, as well as FlexRIC (a flexible RIC), FlexCN (a flexible core controller), and intelligent orchestration tools for RAN and CN domains [104]. Additionally, researchers at Northeastern University have successfully integrated OAI with the OSC RIC, enabling interoperability between open source development and standardized O-RAN implementations. Further details are provided in [105, 106].
@@ II-E5 X5G
II-E7 Insights from Open-Source O-RAN Deployments
Building on the open-source platforms and testbeds introduced earlier, this section synthesizes representative findings from recent evaluations of open 5G deployments.
Our study [119] compares the deployment of commercial and open-source O-gNBs in an industrial environment. The findings indicate that the commercial gNBs offer better coverage, while the open-source gNBs achieve lower and more consistent latency. Despite being less energy-efficient, the open-source gNBs provide greater flexibility, easier configuration, and better maintainability, supported by a vibrant developer community. In contrast, the commercial gNBs operate within a closed system, offering limited scope for customization.
The work in [120] emphasizes how component configurations and interdependency significantly influence E2E performance. Through over-the-air measurements of packet loss and one-way delay under concurrent uplink and downlink transmissions, the authors show that open-source O-RAN-based systems can, under certain conditions, match or surpass conventional solutions in performance, while offering industrial-grade connectivity at lower cost. Interoperability is evaluated in [121], where various open-source RAN and CN combinations are tested, providing critical insights into integration and performance trade-offs.
A performance evaluation presented in [122] highlights the suitability of the srsRAN platform for private 5G network deployments. The study also identifies key limitations, emphasizing that the full potential of 5G remains unrealized—primarily due to constraints associated with general-purpose hardware and the current maturity level of open-source software components. The study [123] employs an OAI-based testbed to evaluate O-RAN performance, with a focus on throughput measurements. The findings expose limitations of the setup and suggest potential underlying causes.
The authors in [124] present a standalone testbed to evaluate downlink performance across two O-RAN implementations: srsRAN and OAI. The results indicate that functional split architectures enhance flexibility and resource efficiency by offloading processing to the O-DU. In [125], the authors propose a cloud-native O-RAN testbed that supports dynamic deployment and O-DU scaling. When central processing unit (CPU) load exceeds 80%, an additional O-DU is instantiated, improving throughput and latency under heavy traffic and demonstrating efficient network resource management.
The demonstration in [126] presents a modular, cloud-native 5G O-RAN, deployed across multiple locations. It enables flexible resource allocation and supports on-demand monitoring applications, facilitating autonomous, and feedback-driven orchestration. The study in [127] evaluates resource scheduling strategies within an O-RAN-compliant 5G network using the ns-3 simulator. By comparing different scheduling mechanisms, the authors demonstrate that the integration of RIC-driven control enhances flexibility and adaptability in scheduling policies.
Further advancing practical implementation, the work in [128] provides a comprehensive guide for developing and deploying O-RAN applications in both simulated and real-world environments. It equips developers with methodologies for architectural evaluation, xApp migration to testbeds, and deployment of key components such as the SMO, Non-RT RIC, and rApps to enable full E2E integration.
@@ III-B9 NG Interface
The NG interface, adopted from 3GPP, connects the O-CU to the 5G core (5GC). It comprises two components: NG-C for the control plane and NG-U for the user planes. NG-C connects the O-CU-CP with the access and mobility management function (AMF), while NG-U connects the O-CU-UP to the user plane function (UPF) [20].
@@ III-C RAN Intelligent Controller
Moreover, the RIC is responsible for configuring network slices, orchestrating NFs, monitoring network performance, and conducting real-time optimizations of RAN resources via open interfaces [158, 160]. As illustrated in Figure 2, the RICis implemented in two distinct forms—Non-RT RIC and Near-RT RIC—each designed to address specific control loop dynamics and latency requirements [149]. The following sections provide a detailed examination of both variants, elaborating on their functionalities, applications, and strategic importance in the broader context of RAN optimization and automation.
@@ III-D The SMO Framework and its Corresponding Interfaces
The architecture of SMO—particularly the Non-RT RIC—offers flexible implementation options, enabling the operators to selectively adopt the desired features. The SMO can integrate with an E2E multi-domain service orchestrator, coordinating domain-specific modules across the RAN, TN, and CN. It facilitates on-demand creation and management of network slice instances (NSIs) over distributed 5G infrastructure [176].
@@ III-D2 O2 Interface
In addition to lifecycle management of O-Cloud infrastructure, the O2 interface supports the orchestration of resource management such as inventory, monitoring, provisioning, and software management. It provides logical services that govern the deployment of O-RAN NFs on cloud resources. The O2 interface is inherently extensible, allowing new features to be incorporated without altering the underlying protocol or management processes. It is vendor-neutral and is unaffected by particular SMO framework and O-Cloud implementations. Through the O1 and/or O2 interfaces, operators can dynamically manage, reconfigure, and upgrade network components hosted within the O-Cloud environment [179].
@@
The 3GPP-defined MFs for the M&O of network slicing include the communication service management function (CSMF), network slice management function (NSMF), network slice subnet management function (NSSMF), and network function management function (NFMF) [38]. These MFs can be integrated within the SMO framework in accordance with the requirements outlined in the 3GPP specifications. Additionally, the O-RAN Alliance has defined both functional and non-functional requirements for network slicing architecture, as detailed in [44]. The requirements specified by both 3GPP and the O-RAN Alliance are critical for the successful realization of network slicing within the O-RAN architecture and for ensuring the effective operation of the MFs. The provision of management services (MnSs) for mobile networks—including network slicing—can be achieved through a set of functional blocks, as illustrated in Figure 3.
@@ III-E3 O-RAN Cloud Platform
The O-RAN cloud platform (O-Cloud) refers to a cloud computing environment composed of physical network infrastructures. It supports the deployment of critical O-RAN components such as the SMO, Near-RT RIC, and E2 nodes, along with the associated software and required M&O services. Each O-Cloud is composed of a group of CPUs, random access memory (RAM), storage, network interface cards (NICs), basic input and output system (BIOS), baseboard management controllers (BMCs), and hardware accelerators. These elements work together to handle computationally intensive tasks across the platform [20].
@@ IV-B3 Service Layer
It is important to recognize that while the requirements for PNFs, cloudified network services, and O-Cloud platforms may vary. However, the logical network function requirements always remain the same [44, 178]. For Example, in the scenario illustrated, a single O-CU-CP instance manages the control plane operations for both network slices, whereas each slice is assigned a dedicated O-CU-UP instance. If the UE is connected to both slices, only one RRC connection is established, handling handover and cell assignments through the shared O-CU-CP. However, each service associated with a different NSI can benefits personalized QoS management and independent flow control through an individual SDAP/PDCP stacks within their respective O-CU-UP [44].
@@ IV-C SMO Framework Deployment Options
The EXT-API offers northbound interoperability for the ONAP platform, serving as an access point for third-party frameworks. Upon receiving a service request, EXT-API responds with a Service Order ID, which can be used to track the status of the service order. Subsequently, the EXT-API activates the service orchestrator (SO) API to initiate the actual service creation process. This action represents progress in establishing uniform external interfaces for automated network slice orchestration [103, 198].
@@ V-C Xhaul Transport Network
V-C1 Xhaul TN Architecture
@@
Underlay networks form the physical infrastructure of a TN, comprising Ethernet switches, routers, dense WDM (DWDM) equipments, and the fiber optic cabling that interconnects these components into a coherent topology. To support an Xhaul TN environment, the packet switched network must handle both L2 and L3 services. Currently, L2 underlay networks predominantly rely on Ethernet, often utilizing virtual LANs (VLANs) for segmentation. Within O-RAN, following two prevalent packet switched underlay technologies are discussed in [202].
@@
V-C2 Xhaul TN Slicing
The packet switched TNs provide a robust framework to support the network slicing. The Xhaul TN infrastructure is designed to support the diverse transport demands. They require tailored solutions for control, management, and user plane interfaces. This includes 5G TN segments, where FH, MH, and BH interfaces require tailored resource allocation [77, 211]. Each interface has specific latency, bandwidth, and traffic demands, which slicing helps to address by enabling precise resource customization [212].
@@ VI-B2 Slow Loop RAN Slice SLA optimization
In the second option the SMO framework updates the slice configuration of the Near-RT RIC or RAN nodes based on instructions from the AI/ML model or control app. After the update, two outcomes are possible: either the Near-RT RIC and the RAN nodes implement the updated configuration, or the Near-RT RIC receives the updated A1 policies, take control of the RAN nodes and provide feedback to the Non-RT RIC.
@@ VII-A Lessons Learned Related to the Architecture of O-RAN
We discussed on several occasions that the O-RAN architecture is composed of key components and interfaces that support its open, intelligent, and modular design. We gained several valuable insights from our study on these components and interfaces. The key takeaways are summarized below.
Fully Disaggregated Cloud-based RAN: The adoption of O-RAN marks a pivotal shift towards fully disaggregated RAN architectures, where key O-gNB functions are modularized into distinct components such as the O-CU, O-DU, and O-RU. The key contribution of O-RAN Alliance is the definition of O-FH interface that splits the DU into O-DU and O-RU. We learned that such a disaggregation can play a significant role in the isolation of O-RAN slices. To further enhance isolation, resource efficiency, and support for cloud-native solutions, it may be beneficial for network operators and vendors to further split the O-CU and O-DU into micro VNFs, such as the virtual network function components (VNFCs) defined by the ETSI ISG NFV.
@@ IX Concluding Remarks and Future Outlook
In conclusion, the exploration of O-RAN illuminates its transformative potential within the context of cellular communications systems. As the wireless industry evolves to meet the demands of 5G, 6G, and beyond, O-RAN emerges as a promising paradigm shift, offering flexibility, interoperability, and cost-efficiency in telecommunications networks deployment and management. Through our comprehensive analysis in this paper, it becomes evident that O-RAN’s disaggregated approach to wireless network elements, enabled by open interfaces, automation, intelligence, and SDN principles, fosters innovation and competition among vendors while reducing vendor lock-in. This approach not only spurs the development of diverse and specialized NFs but also empowers operators to tailor their networks to specific use cases and environments with greater agility and granularity through the deployment of network slicing at both network and management domains. In essence, while the journey towards realizing the full potential of network slicing in O-RAN may be fraught with several challenges, the destination promises a network architecture that is more open, agile, and responsive to the evolving needs of wireless communication in the next decade.